Imagine you borrow $600 from your roommate, agreeing to pay her back $600 plus 7 percent nominal interest in one year. Assume inflation over the life of the contract is expected to be 4.62 percent. What is the total dollar amount you will have to pay her back in a year? What approximate percentage of the interest payment is the result of the real rate of interest?
Answer #1

Amount paid back in 1 year = $600+ 600*7% = $642 Real rate of Interest = (1+7%)/(1+4.62%) -1 = 2.27% Real part of Interest =